# Thermal evolution of dark matter in the early universe from a symplectic glueball model

### v1

@@ 4 Interpretation as a model for dark matter
Further implications of a dark $\operatorname{\mathrm{SU}}(N)$ theory coupled to an axion-like inflaton, resulting in the excitation of the Yang–Mills fields already in the slow-roll inflationary phase (the “minimal warm inflation” scenario [261, 262]), were later analyzed in ref. [93]. In particular, this work studied in detail the cosmological evolution, from inflation to SM nucleosynthesis, of a dark sector that gets populated first, by inflaton decays, and then, in turn, leads to production of particles of the Standard Model through portal interactions [263, 264, 265]. For an $\operatorname{\mathrm{SU}}(N\geq 3)$ theory, the authors of ref. [93] demonstrated that the lightest glueball state with charge-conjugation quantum number $C=-1$ could be a viable dark-matter candidate, with appropriate relic abundance, and with a correct reheating temperature. A key insight leading to these findings is that, while the lightest $C$-even state can couple to the Standard Model Higgs field through a dimension-six operator, the lightest $C$-odd state can only decay via operators of dimension at least equal to eight (for related work, see also refs. [266, 267, 94]). This implies that the decay width of the axial-vector $C=-1$ glueball to a dark $J^{PC}=0^{++}$ glueball and to electroweak bosons is strongly suppressed, thus the lightest $C$-odd state can have a lifetime longer than the age of the universe. This mechanism, whereby the stability of the dark-matter candidate on cosmological time scales is protected by charge conjugation, is very appealing, and, as discussed in ref. [93] leads to very interesting phenomenology, reconciling the requirements for dark matter with those from Standard-Model nucleosynthesis. It should be noted that, unfortunately, this mechanism cannot be at work in our model, which, being based on an $\operatorname{\mathrm{Sp}}(N)$ group, cannot accommodate any $C$-odd state.
@@ 5 Conclusions
While in the past several works have already considered the study of a “dark $\operatorname{\mathrm{SU}}(N)$ theory”, and discussed its potential implications for the evolution of the early universe, much less attention has been devoted to dark-matter models based on confining theories constructed from other non-Abelian gauge groups. This motivated us to consider, in this work, the thermal evolution of a dark-matter model based on the $\operatorname{\mathrm{Sp}}(2)$ theory: the latter shares many qualitative (and, in part, quantitative) features with those based on a special unitary group, but also differs from them for certain aspects. For the purposes of this work, the most interesting differences between $\operatorname{\mathrm{Sp}}(2)$ and $\operatorname{\mathrm{SU}}(N)$ (for $N\geq 3$) Yang–Mills theories are probably the absence of states with a negative quantum number under charge conjugation, and the fact that the center of the group is $\operatorname{\mathbb{Z}}_{2}$. In particular, the non-existence of $C=-1$ glueball states precludes the viability of the scenario recently discussed in ref. [93], which can simultaneously account for correct values of dark matter abundance and for a sufficiently high reheating temperature, and which has the attractive feature of “protecting” the existence of the lightest $C=-1$ state by a discrete symmetry. This mechanism allows the dark matter candidate to survive and be cosmologically stable, even if all other dark glueball states, including the $C=1$ glueballs, decay before primordial nucleosynthesis.
@@
The study of the equation of state of the dark $\operatorname{\mathrm{Sp}}(2)$ theory that we presented in section 3 is entirely non-perturbative, and was carried out in a completely gauge-invariant approach based on first principles, according to the lattice regularization [35]. We stress that this approach provides a mathematically rigorous definition of the theory, which, in contrast to phenomenological models, is free from uncontrolled approximations and is systematically improvable; in the thermodynamic and continuum limits, all discretization artefacts (including, in particular, those related to the explicit breaking of Lorentz–Poincaré invariance by the regularization on a grid) vanish. The motivation for carrying out the present study of the $\operatorname{\mathrm{Sp}}(2)$ equation of state is that, at the temperatures close to $T_{\mbox{\tiny{c}}}$ that we considered, the physics of non-Abelian gauge theories involves non-trivial non-perturbative aspects, which cannot be captured by weak-coupling expansions. While perturbative predictions in high-temperature gauge theories can be improved by combining them with effective field theories [269, 270, 271, 272] constructed via dimensional reduction [273, 274], and this approach is indeed actively pursued in the study of cosmological phase transitions [275], our present study does not rely on any weak- (nor strong-) coupling approximation.
@@
The results for the equilibrium thermodynamic quantities of the $\operatorname{\mathrm{Sp}}(2)$ theory that we presented in section 3 show remarkable similarities with those obtained in $\operatorname{\mathrm{SU}}(N)$ gauge theories [60], and especially with those from the $\operatorname{\mathrm{SU}}(3)$ theory [53, 54]. In both cases, the confinement/deconfinement transition is of first order, with a finite, but not excessively large latent heat: in the symplectic theory, the dimensionless ratio $\frac{L_{\mbox{\tiny{h}}}}{T_{\mbox{\tiny{c}}}^{4}}$ takes a value that is somewhat larger, but very close, to the analogous ratio for the $\operatorname{\mathrm{SU}}(3)$ theory. These close similarities between the theories may be interpreted invoking large-$N$ arguments, and/or the quite similar dimensions of the underlying Lie algebras ($10$ versus $8$).
@@
In view of these results, in section 4 we then discussed some aspects relevant for the interpretation of our theory as a framework for a dark-matter candidate, focusing on its evolution in the early universe. In particular, we concentrated on aspects related to phenomena beyond thermodynamic equilibrium, including some that may be of relevance for the production of gravitational waves. While a systematic lattice study of several physical quantities mentioned therein for the $\operatorname{\mathrm{Sp}}(2)$ theory is beyond the scope of the present work (requiring simulations based on non-trivial generalizations of those that could be used for the equation of state, and demanding a substantial amount of CPU time on supercomputers), we mostly repeated arguments that had already been formulated for dark $\operatorname{\mathrm{SU}}(N)$ theories, trying to highlight analogies and differences relevant for the $\operatorname{\mathrm{Sp}}(2)$ model. All in all, our findings can be summarized by stating that the $\operatorname{\mathrm{Sp}}(2)$ theory considered in this work is a viable and potentially interesting dark-matter model.
@@
While we tried to make the numerical study as complete as was feasible, and our discussion in this paper as exhaustive as possible, there are several directions in which the present work could be generalized. Our lattice determination of equilibrium thermodynamic quantities could be refined with simulations on larger lattices (to improve the extrapolations to the thermodynamic limit) and at finer spacings (to improve the extrapolations to the continuum limit); in addition, the temperature range that was probed could be extended. These extensions of the numerical calculations presented herein would reduce the systematic and statistical uncertainties of our results for the equation of state of the $\operatorname{\mathrm{Sp}}(2)$ theory, but would require a non-negligible amount of computational resources. Even though the lattice study of systems featuring a first-order transition involves its own challenges, the technology to tackle them is now available [139]. A different type of generalization of this work would consist in trying to compute quantities beyond equilibrium thermodynamics. Examples of such calculations have already been performed for the $\operatorname{\mathrm{SU}}(3)$ gauge theory, and their extension to a symplectic Yang–Mills theory would be new. It would also be interesting to repeat calculations such as those that we discussed in section 4 using parameters that can be extracted (or, at least, estimated) from dedicated lattice simulations of the $\operatorname{\mathrm{Sp}}(2)$ theory. Finally, other possible generalizations of this work could be to investigate other $\operatorname{\mathrm{Sp}}(N>2)$ theories, or theories based on different gauge groups, and/or coupling the theory with dynamical matter fields. We leave all of these interesting possibilities for future research.

### v2

@@ 4 Interpretation as a model for dark matter and gravitational-wave production
Further implications of a dark $\operatorname{\mathrm{SU}}(N)$ theory coupled to an axion-like inflaton, resulting in the excitation of the Yang–Mills fields already in the slow-roll inflationary phase (the “minimal warm inflation” scenario [357, 358]), were later analyzed in ref. [106]. In particular, this work studied in detail the cosmological evolution, from inflation to SM nucleosynthesis, of a dark sector that gets populated first by inflaton decays, and then, in turn, leads to production of particles of the Standard Model through portal interactions [359, 360, 361]. For an $\operatorname{\mathrm{SU}}(N\geq 3)$ theory, the authors of ref. [106] demonstrated that the lightest glueball state with charge-conjugation quantum number C=1C=-1 could be a viable dark-matter candidate, with appropriate relic abundance, and with a correct reheating temperature. A key insight leading to these findings is that, while the lightest $C$-even state can couple to the Standard Model Higgs field through a dimension-six operator, the lightest $C$-odd state can only decay via operators of dimension at least equal to eight (for related work, see also refs. [362, 363, 107]). This implies that the decay width of the axial-vector C=1C=-1 glueball to a dark $J^{PC}=0^{++}$ glueball and to electroweak bosons is strongly suppressed, thus the lightest $C$-odd state can have a lifetime longer than the age of the universe. This mechanism, whereby the stability of the dark-matter candidate on cosmological time scales is protected by charge conjugation, is very appealing, and, as discussed in ref. [106], leads to very interesting phenomenology, reconciling the requirements for dark matter with those from Standard-Model nucleosynthesis. It should be noted that, unfortunately, this mechanism cannot be at work in our model, which, being based on an $\operatorname{\mathrm{Sp}}(N)$ group, cannot accommodate any $C$-odd state.
A different situation can be realized under the hypothesis that the dark sector is not reheated by the decay of the inflaton, and is instead populated only by gravitational interactions with the visible sector at very high temperatures, deep in the deconfined phase [364, 365]. In particular, it was shown in ref. [362, section 3] that in this scenario the problem of dark-matter overproduction can be evaded.
In general, understanding whether the parameter space of a certain theory is consistent with observational constraints (in particular those from Big-Bang nucleosynthesis and from the cosmic microwave background) is a central question of any particle dark matter model. For the gauge theory that we considered, we can distinguish the relevant parameters into two groups:
  • parameters related to general, “static” or “equilibrium” properties of the theory: these include, in particular, the masses of physical states in the confining phase, the deconfinement temperature, the latent heat, etc.;
  • parameters related to “dynamical” phenomena: examples of these include, for instance, glueball decay and/or annihilation rates.
While the lattice calculations that we presented in this work allow one to determine unambiguously, in a gauge-invariant, fully non-perturbative and systematically improvable way all of the parameters within the first group (except one, which has to be used to set the scale), they do not give us direct access to those in the second group, for which we have to rely on general arguments or on crude tentative estimates.
As we remarked above, dark matter contributes to the total energy density of the universe, and thus may affect, in particular, the primordial deuterium and helium-$4$ yields; at the same time, it may also affect the cosmic microwave background, since it can modify the tail of temperature-polarization spectra. It should be noted that relativistic particle species dominate the early‑Universe expansion, as their contribution to the energy density is proportional to the fourth power of the temperature. This case is relevant in the high-temperature regime of the $\operatorname{\mathrm{Sp}}(2)$ model that we discussed here, in which our lattice results for the equation of state show that the energy density scales approximately like the fourth power of the temperature (up to a further, mild dependence on the temperature, which accounts for the fact that, for temperatures not much higher than the deconfinement temperature, the “gluon-like” quasiparticles in the deconfined phase are not completely free; rather, the coupling describing their mutual interactions is expected to scale with the inverse of the logarithm of the temperature). Conversely, the contribution to the energy density from massive species is exponentially suppressed: when the temperature drops below the particle mass, the particle becomes non-relativistic, and the kinetic energy can be neglected with respect to the rest-mass energy: this implies that the particle’s contribution to the energy density becomes proportional to the exponential of the ratio of the particle mass to the temperature, times a power of the temperature—see also the expression for the number density in eq. (4.26). As we discussed at the beginning of section 3, for the $\operatorname{\mathrm{Sp}}(2)$ gauge theory the masses of all physical states in the spectrum are significantly larger than the critical temperature at which the theory enters its confining phase and develops a mass gap: this implies that, as soon as the temperature of the theory drops below $T_{\mbox{\tiny{c}}}$, the mass-to-temperature ratio is quite large (larger than $5$ for the lightest glueball species), and thus the energy-density contribution from the dark sector becomes strongly suppressed.
A concrete example of a confining dark sector, in which the problem of dark-matter overproduction can be evaded, was presented in ref. [362]: while the focus of that work was on a dark $\operatorname{\mathrm{SU}}(3)$ Yang–Mills theory, we note that most of their arguments apply in the case of a dark $\operatorname{\mathrm{Sp}}(2)$ gauge theory, too, up to small quantitative changes: for example, the central charge of the theory in the plasma phase is simply rescaled by a factor equal to the ratio of the number of gluon degrees of freedom in the two theories, i.e., $10/8=1.25$, and similar changes can be applied for the quantities relevant for the glueball phase, like the value of the mass of the lightest glueball and the latent heat in the appropriate units of the deconfinement temperature. In particular, it is interesting to discuss two different situations, namely the case in which the number density of glueballs produced at the confining transition is not significantly modified by subsequent number-changing processes, and the case in which, instead, $3\to 2$ processes alter it: since in this case dark matter particles gain energy “by eating each other”, this latter scenario is often referred to as “cannibalism” [29].
If dark glueball interactions are sufficiently weak and one can neglect processes that change the glueball number density, the latter can be estimated from energy conservation at the phase transition (under the assumption used above, that the nucleation temperature is very close to the critical temperature). Note that, as we mentioned above, the assumption of negligible glueball interactions in a gauge theory without elementary matter fields in the fundamental representation is at least qualitatively supported by large-$N$ arguments [340], and the most recent lattice results suggest that even the purely geometric cross-section of glueballs may be small [366]. Comparison with experimental results from the strong-interaction sector of the Standard Model, on the other hand, is of little guidance, since in QCD glueballs are affected by significant mixing with flavor-singlet mesons, and, in fact, even the proper identification of glueball states in QCD remains somewhat elusive [367, 368, 369, 370]. In any case, if one neglects glueball interactions, then, following the discussion in ref. [362, section 3] and noting how the parameters of the $\operatorname{\mathrm{Sp}}(2)$ theory are related to those of the $\operatorname{\mathrm{SU}}(3)$ theory, one obtains a tentative estimate for the yield:
$Y\simeq 2\cdot 10^{-3}\left(\frac{T_{\mbox{\tiny{reh}}}}{m_{\mbox{\tiny{Planck}}}}\right)^{9/4},$ (4.32)
where $T_{\mbox{\tiny{reh}}}$ denotes the reheating temperature, and a corresponding crude estimate for the dark matter abundance:
$\frac{\Omega h^{2}}{0.12}\simeq\frac{m_{0^{+}}}{10~\mbox{GeV}}\left(\frac{T_{\mbox{\tiny{reh}}}}{10^{15}~\mbox{GeV}}\right)^{9/4}.$ (4.33)
It should be stressed that, in view of the inherent uncertainties on the assumption of negligible glueball interactions, eq. (4.32) and eq. (4.33) should be considered as order-of-magnitude estimates only.
On the other hand, if the glueball interactions are to be taken into account, one can estimate the yield by considering the effect of $3\to 2$ interactions, which are expected to be those responsible for the leading glueball number density changing processes. Note that, parametrically, the cross-section associated with these interactions is inversely proportional to the fifth power of the glueball mass, hence it is strongly suppressed if the dynamically generated scale of the dark sector is large. Nevertheless, $3\to 2$ processes can be sufficiently frequent (as compared to the expansion rate of the universe) if the coefficient encoding the quantitative magnitude of these interactions and the ratio of the energy density of the dark sector to the one of the visible sector are not too small, and the glueball mass is not too large. In this case, cannibalism allows glueballs to remain in thermal equilibrium, and the temperature of the dark sector varies slowly, with the logarithm of the cosmological scale factor [29, 83], and, as discussed in detail in ref. [362], the dark matter equilibrium yield is then linear in the temperature of the dark sector. As the temperature continues to decrease, the rate of $3\to 2$ processes eventually drops below the Hubble rate, at which point cannibalism stops, leaving a dark matter yield directly related to the ratio of the entropy densities of the dark and visible sectors. In order to make precise quantitative predictions about this type of dynamics, it would be crucial to know exactly the value of the cross-section of $3\to 2$ interactions. Unfortunately, however, studying this type of processes on the lattice is particularly challenging [371, 372, 373]. At best, we can only make semiqualitative arguments, and/or rely on the expectation that estimates obtained in the case of a dark $\operatorname{\mathrm{SU}}(3)$ theory could approximately hold also for the $\operatorname{\mathrm{Sp}}(2)$ model. In particular, for the $\operatorname{\mathrm{SU}}(3)$ theory, the matrix elements associated to the decay of the lightest glueball were studied in refs. [126, 128]; the relative uncertainties affecting this type of computations are in the ballpark of $20\%$, i.e., significantly larger than those of the glueball spectrum. We expect that the systematic uncertainty in the assumption that the decay matrix elements of the $\operatorname{\mathrm{Sp}}(2)$ theory take values similar to those of the $\operatorname{\mathrm{SU}}(3)$ does not exceed such uncertainties.
In short, the similarities between the model that we considered herein and a dark $\operatorname{\mathrm{SU}}(3)$ Yang–Mills theory suggest that the conclusions reached in ref. [362] are relevant for our model, too, and that, for example, a plot of the viable values for the confining temperature of gravitationally produced $\operatorname{\mathrm{Sp}}(2)$ glueball dark matter as a function of the reheating temperature of the visible sector would be quantitatively very similar to the one shown in ref. [362, figure 1].
A different, and more involved, discussion would be necessary, instead, if one were to explicitly consider also the variety of scenarios by which dark glueballs can eventually decay to the Standard Model via connector operators. For a dark $\operatorname{\mathrm{SU}}(3)$ Yang–Mills theory, these possibilities were discussed in detail in ref. [83], considering, in particular, operators of different dimensions (and with broken or unbroken discrete charge-conjugation symmetry, which is not present in our model). In particular, if the lifetimes of dark glueballs are sufficiently short, their decays can induce changes in the primordial yields of light elements produced through Big Bang nucleosynthesis [347, 374], they can modify the cosmic microwave background [345, 375], and alter the spectrum of cosmic rays [344]. In the absence of an accurate, quantitative estimate for the parameters relevant for such decays in the $\operatorname{\mathrm{Sp}}(2)$ model, however, we do not make any comments on this possibility, except that, once again, we expect the overall picture to be roughly similar to the $\operatorname{\mathrm{SU}}(3)$ case, except for the absence of states with negative charge-conjugation quantum number.
To summarize, while in the present work we addressed the explicit lattice calculation of a number of physical quantities for the $\operatorname{\mathrm{Sp}}(2)$ theory, and could therefore derive robust predictions for the equation of state and for (at least certain aspects related to) the gravitational-wave production at the confining transition of this dark Yang–Mills model, for observables that depend on physical quantities that were not addressed in the present lattice calculation, we necessarily have to rely on qualitative or semi-quantitative estimates only, based on theoretical arguments (the large-$N$ limit and effective field theory) or from some empirical numerical evidence (including, in particular, the similar values of quantities computed from lattice simulations of the $\operatorname{\mathrm{Sp}}(2)$ theory and of the $\operatorname{\mathrm{SU}}(3)$ theory). While this implies that, as we discussed above, some of the expectations derived from the $\operatorname{\mathrm{Sp}}(2)$ model at present remain at the level of educated guesses, and/or are meant to be more qualitative than quantitative, in principle the analysis could still be refined through a sequence of dedicated studies in a longer-term plan. Examples of quantities for which this would be in principle possible include, for instance, the matrix elements associated with glueball decays, for which one could repeat a study similar to those that were carried out for the pure-glue sector of QCD in refs. [126, 128]. Even this small step, however, would require significant computational time and resources, and we postpone it to the future.
@@ 5 Conclusions
While in the past several works have already considered the study of a “dark $\operatorname{\mathrm{SU}}(N)$ theory”, and discussed its potential implications for the evolution of the early universe, much less attention has been devoted to dark-matter models based on confining theories constructed from other non-Abelian gauge groups. This motivated us to consider, in this work, the thermal evolution of a dark-matter model based on the $\operatorname{\mathrm{Sp}}(2)$ theory: the latter shares many qualitative (and, in part, quantitative) features with those based on a special unitary group, but also differs from them under some aspects. For the purposes of this work, the most interesting differences between $\operatorname{\mathrm{Sp}}(2)$ and $\operatorname{\mathrm{SU}}(N)$ (for $N\geq 3$) Yang–Mills theories are probably the absence of states with a negative quantum number under charge conjugation, and the fact that the center of the group is $\operatorname{\mathbb{Z}}_{2}$. In particular, the non-existence of C=1C=-1 glueball states precludes the viability of the scenario discussed in ref. [106], which can simultaneously account for correct values of dark matter abundance and for a sufficiently high reheating temperature, and which has the attractive feature of “protecting” the existence of the lightest C=1C=-1 state by a discrete symmetry. This mechanism allows the dark matter candidate to survive and be cosmologically stable, even if all other dark glueball states, including the $C=1$ glueballs, decay before primordial nucleosynthesis.
@@
The study of the equation of state of the dark $\operatorname{\mathrm{Sp}}(2)$ theory that we presented in section 3 is entirely non-perturbative, and was carried out in a completely gauge-invariant approach based on first principles, in the lattice regularization [36]. We stress that this approach provides a mathematically rigorous definition of the theory, which, in contrast to phenomenological models, is free from uncontrolled approximations and is systematically improvable; in the thermodynamic and continuum limits, all discretization artifacts (including, in particular, those related to the explicit breaking of Lorentz–Poincaré invariance by the regularization on a grid) vanish. The motivation for carrying out the present study of the $\operatorname{\mathrm{Sp}}(2)$ equation of state on the lattice is that, at the temperatures close to $T_{\mbox{\tiny{c}}}$ that we considered, the physics of non-Abelian gauge theories involves non-trivial non-perturbative aspects, which cannot be captured by weak-coupling expansions. While perturbative predictions in high-temperature gauge theories can be improved by combining them with effective field theories [377, 378, 379, 380] constructed via dimensional reduction [381, 382], and this approach is indeed actively pursued in the study of cosmological phase transitions [300], our present study does not rely on any weak- (nor strong-) coupling approximation.
@@
The results for the equilibrium thermodynamic quantities of the $\operatorname{\mathrm{Sp}}(2)$ theory that we presented in section 3 show remarkable similarities with those obtained in $\operatorname{\mathrm{SU}}(N)$ gauge theories [70], and especially with those from the $\operatorname{\mathrm{SU}}(3)$ theory [63, 64]. In both cases, the confinement/deconfinement transition is of first order, with a finite, but not excessively large latent heat: in the symplectic theory, the dimensionless ratio $L_{\mbox{\tiny{h}}}/T_{\mbox{\tiny{c}}}^{4}$ takes a value that is larger than, but of the same order of, the one it has in the $\operatorname{\mathrm{SU}}(3)$ theory. These similarities between the theories may be interpreted invoking large-$N$ arguments, and/or the quite similar dimensions of the underlying Lie algebras ($10$ versus $8$).
In section 4 we then discussed some topics relevant for the interpretation of our model as a framework for a dark-matter candidate, focusing on its evolution in the early universe. In particular, we concentrated on aspects related to phenomena beyond thermodynamic equilibrium, including some that may be of relevance for the production of gravitational waves. It is important to remark that gravitational waves generated by cosmological processes in the early universe will appear today as randomly distributed in all directions, with many unresolved sources. As a matter of fact, the number of Hubble patches at a given redshift $z$ that fit the current universe scales as $(1+z)^{3}$: this is a huge number for gravitational waves generated at large $z$, so the superposition of all incoming unresolved primordial gravitational waves results into a stochastic gravitational-wave background [383], and, depending on the temperature of the universe at the time of the gravitational wave production, the corresponding peak frequencies can be in the range accessible to different detectors. As we discussed, gravitational waves can be sourced by bubble collisions, by sound waves, or by hydrodynamic fluctuations of a plasma of particles at high temperature. The latter mechanism has been studied in detail and extended, in an approach directly based on the “microscopic” formulation of the Standard Model, in ref. [44], where it was found that the energy density carried by thermally produced gravitational radiation is maximal for gravitational waves with frequencies of the order of $10^{11}$ Hz. Such frequencies are very far from the optimal operational range of present experiments, and, while this may change with future detectors [384, 385, 386, 387], reaching the necessary sensitivity would likely remain a challenge. As was pointed out in ref. [388], however, the situation may be quite different in the presence of a new dark sector, like the one that we considered in the present work.
@@
Albeit a more systematic lattice study of several physical quantities mentioned in section 4 for the $\operatorname{\mathrm{Sp}}(2)$ theory remains beyond the scope of the present work (as it would require simulations based on non-trivial generalizations of those that we performed for the equation of state, and demanding a substantial amount of CPU time on supercomputers), so, as we already mentioned, the results presented in section 4 should be interpreted as benchmark examples, we remark that our approach followed the main arguments that had already been put forward for dark $\operatorname{\mathrm{SU}}(N)$ theories, trying to highlight analogies and differences relevant for the $\operatorname{\mathrm{Sp}}(2)$ model. Our results can also be compared with those obtained using different techniques to study strongly coupled gauge theories, like the gauge/gravity duality [389, 390, 391]: examples of such studies, with a particular focus on the spectrum of gravitational waves, include those reported in refs. [90, 392]. All in all, our findings can be summarized by stating that the $\operatorname{\mathrm{Sp}}(2)$ theory considered in this work is a viable and potentially interesting dark-matter model, and that, depending on the value of the energy scale characterizing the model, the gravitational-wave spectrum produced at its first-order deconfinement/confinement phase transition could have a potentially observable signature, which could reveal important information on the early stages of the universe evolution.
@@
While we tried to make our numerical study as complete as was feasible, and the discussion in this paper as exhaustive as possible, there are several directions in which the present work could be generalized. Our lattice determination of equilibrium thermodynamic quantities could be refined with simulations on larger lattices (to improve the extrapolations to the thermodynamic limit) and at finer spacings (to improve the extrapolations to the continuum limit); in addition, the temperature range that was probed could be expanded. These extensions of the numerical calculations presented herein would reduce the systematic and statistical uncertainties of our results for the equation of state of the $\operatorname{\mathrm{Sp}}(2)$ theory, but would require a non-negligible amount of computational resources. Even though the lattice study of systems featuring a first-order transition involves its own challenges, the technology to tackle them is now available [153]. A different type of generalization of this work would consist in trying to carry out a non-perturbative computation of quantities beyond equilibrium thermodynamics, like the bubble nucleation rate [266] or the interface tension between the confining and deconfined phases [393]. While such calculations have already been performed for $\operatorname{\mathrm{SU}}(N)$ gauge theories (for recent work, see refs. [394, 395, 396]), their extension to a symplectic Yang–Mills theory would be new. Then, it would also be interesting to repeat calculations such as those that we discussed in section 4 using the parameters that can be extracted (or, at least, estimated) from dedicated lattice simulations of the $\operatorname{\mathrm{Sp}}(2)$ theory. Finally, other possible generalizations of this work could be to investigate other $\operatorname{\mathrm{Sp}}(N>2)$ theories, or theories based on different gauge groups, and/or coupling the theory with dynamical matter fields. We leave all of these interesting possibilities for future research.